Find the inverse of the one to one function : f(x)= 2/7x + 3

Answer by stanbon(48535)

I'm going to assume your problem is as follows:
y = (2/7)x + 3
To form the inverse of this function:
1. Interchange "x" and "y" as follows
x = (2/7)y + 3
2. Now solve this equation for "y" as follows
Subtract "3" from both sides of the equation to get
x - 3 = (2/7)y
Multiply both sides by (7/2) to get
(7/2)(x-3) = y
Rewrite as follows
y = (7/2)(x-3)
Rewrite again as follows
y = (7/2)x - (21/3)
This is the inverse in the form y=mx+b
Good luck
Cheers,
Stan H.